# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

### v1

@@ 3.2 Data Quality Evaluation
To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP radford2021learningtransferablevisualmodels . This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks. For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective. The human-centric template is defined as: A human is [action]”; the robot-centric template is defined as: A robotic arm is [action]” . Here, [action] is a natural language phrase describing the high-level behavior, such as holding a bottle” or “brushing paint on a wall.”
@@
We use the CLIP ViT-B/32 model to compute cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt. This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description. Similarly, we compute the similarity between the original human frame and the human-centric prompt. This setup allows us to directly compare semantic alignment before and after augmentation under their respective modalities. For each frame, only the action-relevant region is retained, and background context is preserved via inpainting using LaMa. We average similarity scores over multiple samples per action class to mitigate variance and isolate the impact of robotic replacement. All textual prompts are manually curated to maintain consistency across frames and embodiments.
This evaluation protocol provides a lightweight, scalable, and interpretable metric for measuring the effectiveness of visual hand-to-robot transformations in preserving task semantics. Detailed results and further analyses are presented in Appendix D.
3.3 H2R Dataset Construction
We construct four large-scale robot-centric video datasets using the proposed H2R augmentation pipeline, each combining a specific robot embodiment with a human egocentric video dataset. These datasets are designed to support policy learning, visual representation learning, and embodiment generalization.
Applying H2R to SSv2. We begin with the SSv2 dataset, which contains 220,847 video clips of human actions with everyday objects, designed to help models understand fine-grained hand gestures. From the official training split, we select 62,500 videos that cover a wide variety of manipulation categories. For each video, we uniformly sample 16 keyframes, resulting in exactly 1,000,000 frames. Each sampled frame is passed through the H2R pipeline to replace the human hand with a simulated robotic arm. For the UR5 robot, we construct three distinct datasets with different end-effector configurations: H2R-UR5-SSv2-1M-Gripper uses a standard two-finger parallel gripper; H2R-UR5-SSv2-1M-Leaphand employs a four-finger anthropomorphic Leaphand; and H2R-UR5-SSv2-1M-Mix combines both embodiments at the frame level to increase visual diversity and generalization. For the Franka robot, which is configured only with a two-finger gripper, we construct H2R-Franka-SSv2-1M following the same augmentation strategy.
Applying H2R to Ego4D. To expand the applicability of our dataset to unconstrained, real-world human activity, we apply the same augmentation pipeline to the Ego4D dataset, which features long-form, first-person videos of natural daily behavior across diverse environments. Following the action-centric preprocessing strategy introduced by MPI zeng2024learningmanipulationpredictinginteraction , we extract 117,624 action clips from 2,486 Ego4D videos, each consisting of three keyframes. These are further processed to produce H2R-UR5-Ego4D-1M-Gripper, H2R-UR5-Ego4D-1M-Leaphand, and H2R-UR5-Ego4D-1M-Mix, using the same embodiment configuration logic as in the SSv2 case. For Franka, we similarly construct H2R-Franka-Ego4D-1M using the gripper configuration.
Each dataset contains approximately one million augmented images. In addition to the RGB images, we provide comprehensive frame-level metadata to support downstream use. Each entry includes: (1) the original label (from SSv2 or Ego4D), (2) the 3D hand keypoints, camera intrinsics, and extrinsics estimated by HaMeR, (3) the inpainted background image generated by LaMa, and (4) the final robot joint configurations and camera pose alignment used in the simulator for rendering. All robotic arms are rendered using the SAPIEN simulator, with camera viewpoint precisely matched to the original egocentric perspective using a two-step transformation between the HaMeR-estimated camera pose and the robot-world coordinate frame. Each dataset is released in a modular file structure, with paired directories for images, masks, and JSON metadata files. This organization ensures flexible access for both visual pre-training pipelines and embodied policy learning frameworks.
@@ 4 Simulation Experiment
4.1 Experiment Setup
Figure 4: Simulation Benchmark Overview. Visualization of the seven simulation tasks used for evaluation, including three from Robomimic (MoveCan, Square, Lift), one PushT task from Diffusion Policy, and three from RLBench (Close Box, Close Laptop Lid, Toilet Seat Down).
Encoder Pre-training. We adopt the MAE he2022mae ; xiao2022mvp and R3M nair2022r3m frameworks for pre-training, each employing a Vision Transformer (ViT) Base dosovitskiy2021imageworth16x16words model as the visual encoder. The training dataset used is either Ego4D (117K clips) or H2R-UR5-SSv2-1M-Mix (1M H2R-augmented images). Both MAE and R3M are pre-trained with 8$\times$×A800 GPUs. MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate. For policy training, we present the set of parameters that necessitate policy learning in Appendix  C.
Simulation Benchmark. For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning. Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic mandlekar2021matters , RLbench james2019rlbench , and Diffusion Policy chi2023diffusionpolicy . In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects. For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down. We also use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location. An overview of these simulation tasks is shown in Figure 4.
Table 1: Simulation Benchmark Results. Success rates (%) across imitation learning tasks in Robomimic, PushT, and RLBench environments, evaluated with MAE and R3M encoders before and after applying H2R. Blue represents an increase in task success rate, while Red represents a decrease. All subsequent tables follow the same rule.
Robomimic Diffusion Policy RLBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (SSv2) 54 25.5 94.5 58 59.2 0 10 0 3.3 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5 (+5%) 15 (+5.0%) 20 (+20.0%) 13.3 (+10%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5 (+5.0%) 20 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%)
@@ 4.2 Simulation Results
Table 1 shows that the encoders trained on H2R-UR5-SSv2-1M-Mix dataset significantly outperform those trained on the original Ego4D data across all simulation tasks. Specifically, for Robomimic tasks, the MAE encoder achieves an average success rate improvement of 10.2%, while the R3M encoder improves by 6.3%. In particular, the MoveCan task exhibits a substantial 25.5% increase for MAE trained on the H2R-1M dataset. For the PushT task, H2R leads to success rate gains of 5.3% (MAE) and 7.0% (R3M), further confirming the generalization benefits brought by our method. Moreover, in the RLBench benchmark, H2R consistently enhances performance, achieving an average improvement of 10.0% for MAE and 5.0% for R3M.
These results demonstrate that H2R-1M dataset delivers superior visual representations for imitation learning compared to large-scale human video datasets like Ego4D. H2R  achieve an improvement in the pre-training performance of the vision encoder on imitation learning by bridging the visual gap between human hands and robotic arms.
4.3 Performance on Other Datasets
@@
Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D)) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.5 (-0.1%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
In addition to the H2R-UR5-SSv2-1M-Mix dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the H2R-UR5-Ego4D-1M-Mix dataset. The experimental results in the simulator are shown in Table 2, which denotes that H2R remains highly effective when applied to the Ego4D dataset.
@@ 5 Real World Experiment
Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 21.7 43.3 (+21.7%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%)
5.1 Experiment setup
Robot Setups. We validate the effectiveness of H2R in real-world manipulation tasks using a UR5 robotic arm with two different end effectors: Gripper RobotiqAdaptiveGrippers and Leaphand shaw2023leaphandlowcostefficient . Realsense IntelRealSenseL515 is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model. Real-world setups are shown in Figure 6.
Figure 5: Visualization of Real-world Manipulation Tasks. The left columns show Gripper tasks and the right columns show Leaphand tasks. Each task is illustrated with six frames, demonstrating the progression from the initial state to the completion of the manipulation.
Real-world Tasks. We set up six tasks for gripper manipulation and dexterous manipulation (Figure 5). For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed. For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened.
Data Collection. We collect expert demonstrations through human teleoperation. For gripper manipulation, we use keyboard-based teleoperation. For dexterous manipulation, we teleoperate Leaphand using a vision-based retargeting system. Due to differences in task complexity, robot embodiment, and teleoperation methods, we adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 4.
@@ 5.1 Experiment setup
Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800
Policy Training. For policy training, we select the Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.1. DP and ACT are trained for 300 epochs for each gripper manipulation task and 3000 epochs for each dexterous manipulation task.
@@
Gripper Setup
Leaphand Setup
Figure 6: Real-world Robot Setups. Experimental setups using a UR5 robot with two different end effectors: a parallel-jaw gripper (left) and a dexterous Leaphand hand (right).

### v3

@@ 3.2 Data Quality Evaluation
To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP [radford2021learningtransferablevisualmodels]. This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks. For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective. The human-centric template is defined as: A human is [action]; the robot-centric template is defined as: A robotic arm is [action]. Here, [action] is a natural language phrase describing the high-level behavior, such as ”holding a bottle” or ”brushing paint on a wall”.
@@
We use the CLIP ViT-B/32 model to compute the cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt. This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description. Similarly, we compute the similarity between the original human frame and the human-centric prompt. This setup allows us to directly compare semantic alignment before and after augmentation under their respective modalities. For each frame, only the action-relevant region is retained, and background context is preserved via inpainting using LaMa. We average similarity scores over multiple samples per action class to mitigate variance and isolate the impact of robotic replacement. All textual prompts are manually curated to maintain consistency across frames and embodiments.
Figure 2 contains six pairs of subfigures, each showing a human action and its corresponding augmented robotic action. The six action pairs in Figure 2 (left to right) are described as follows: (1) Holding a yellow measuring tape at a construction site; (2) Holding a bottle; (3) Painting on the wall with a brush; (4) Rushing paint on a wall; (5) Turning on the kitchen faucet; (6) Preparing painting materials on a table.
The CLIP similarity scores are reported in Table 1. Across all samples, H2R-augmented images consistently achieve higher similarity with robot-centric action descriptions than original images do with human-centric ones. This confirms that H2R augmentation effectively improves visual-semantic alignment between human demonstrations and robotic execution.
Img-1 Img-2 Img-3 Img-4 Img-5 Img-6
Ori 30.6 23.7 31.2 31.7 27.5 28.7
Aug 32.6 28.9 32.3 32.0 28.7 31.0
This evaluation provides a lightweight, scalable, and interpretable metric for assessing how well visual hand-to-robot transformations preserve task semantics. To quantify data quality, we sample 1,000 image pairs (original and H2R-augmented). For each original image, we use Qwen-2.5-VL [bai2025qwen25vltechnicalreport] to generate a verb-noun action description (e.g., “washing bowl”). We then form two prompts: “A human is [action]” and “A robot arm is [action]”. Using CLIP ViT-B/32, we compute the cosine similarity between: (1) the human prompt and the original image, and (2) the robot prompt and the H2R-augmented image. We report the average similarity over all 1,000 samples. The raw images achieve an average score of 28.01, while H2R-augmented images reach 29.83. This demonstrates that H2R-augmented images exhibit stronger semantic alignment with robot-centric action descriptions, confirming improved visual-semantic coherence for robotic manipulation. The results validate our data curation pipeline as effective for generating high-quality, robot-relevant visual data at scale.
@@ 4 Simulation Experiment
Encoder Pre-training. We adopt the MAE [he2022mae, xiao2022mvp] and R3M [nair2022r3m] frameworks for pre-training, each employing a Vision Transformer (ViT) Base [dosovitskiy2021imageworth16x16words] model as the visual encoder. The training dataset used is a subset of Ego4D (117K clips) or SSv2 (1M images). Both MAE and R3M are pre-trained with $\timesA8$00 GPUs. MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate.
Policy Training. We train for 200 steps and report the mean success rate for Robomimic tasks. For tasks in RLBench [james2019rlbench], we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization. For the PushT task, we train the Diffusion Policy model for 200 epochs and report the success rate in the simulation environment. For VC-1 Cortex benchmark, we train policies using behavior cloning composed of a 3-layer MLP for 100 epochs. The training hyperparameters used in this work are identical to those described in the original paper.
Simulation Benchmark. For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning. Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic [mandlekar2021matters], RLbench [james2019rlbench], and Diffusion Policy [chi2023diffusionpolicy]. In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects. For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down. We use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location. We also use three MetaWorld tasks: Assembly, ButtonPress and Hammer from VC-1 Cortex benchmark for evaluation. An overview of these simulation tasks is shown in Appendix.
Table 2: Simulation Benchmark Results. Success rates (%) across imitation learning tasks in Robomimic, PushT, RLBench and CortexBench, evaluated with MAE and R3M encoders before and after applying H2R. Blue represents an increase in task success rate, while Red represents a decrease. All subsequent tables follow the same rule.
Robomimic Diffusion Policy RLBench CortexBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average Assembly ButtonPress Hammer Average MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) 88.0(+4.0%) 88.0(+8.0%) 100.0 (+4.0%) 92.0(+5.3%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 76.0 56.0 88 73.3 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5.0 (+5.0%) 20.0 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%) 68.0 (-8.0%) 60.0 (+4.0%) 96.0 (+8.0%) 74.7 (+1.3%)
@@ 4.1 Simulation Results
Table 2 shows encoders trained on augmented data H2R-SSv2 significantly outperform those on original SSv2 across all simulation tasks. For Robomimic tasks, MAE achieves a 10.2% average success rate improvement and R3M 6.3%, with MAE seeing a substantial 25.5% gain on MoveCan. On PushT, H2R boosts MAE by 5.3% and R3M by 7.0%. In RLBench, MAE improves its performance by 10.0% on average and R3M by 5.0%. In CortexBench, H2R improves MAE by 5.3% and R3M by 1.3%. These results confirm that the H2R-augmented dataset delivers superior visual representations for imitation learning to those learned from large-scale human video datasets like SSv2, as H2R  bridges the visual gap between human hands and robotic arms to enhance vision encoder pre-training performance.
@@
Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.6 (+0.0%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
In addition to SSv2 dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the Ego4D dataset. The experimental results in the simulator are shown in Table 3, which indicates that H2R remains highly effective when applied to the Ego4D dataset.
4.2 Pretraining Model on Robotic Datasets
Table 4: Comparison with model pretrained on robotic datasets. All visual encoder weights are initialized using the R3M pretraining method, with the “R3M-H2R” variant pretrained on the H2R-augmented SSv2 dataset.
Robomimic MoveCan Square Lift Average R3M 59.5 20.5 85.0 55.0 R3M-H2R 61.5 37.5 85.0 61.3 R3M-DROID 54.0 22.0 96.0 56.7
To further evaluate how H2R compares to models pretrained directly on robotic datasets, we refer to a recent study [jiang2024robotspretrainrobotsmanipulationcentric], which conducted a systematic comparison between visual representations pretrained on large-scale egocentric human video datasets and those trained on robotic datasets such as DROID [khazatsky2024droidlargescaleinthewildrobot]. The experimental results, summarized in Table 4, show that models pretrained on robotic datasets often provide limited or even degraded performance on downstream manipulation Robomimic benchmarks. Moreover, despite being smaller and simulated rather than real-robot data, the H2R dataset outperforms the DROID dataset in pre-training effectiveness, indicating its strong potential for robotic model pre-training and reflecting its high data quality.
4.3 Effects of Demonstration Density
The relatively low success rates observed in Table 2 are primarily due to our use of keypoint frames in RLBench tasks. This approach, widely adopted in the RLBench benchmark [goyal2023rvt, jia2024lift3dfoundationpolicylifting, shridhar2023peract], accelerates both training and evaluation by reducing redundant frames. However, selecting only keyframes inevitably decreases the density of demonstration data. Since the diffusion policy employed in our imitation learning framework benefits from dense and temporally continuous data due to its long observation horizon, sparse sampling may limit its performance.
To further investigate demonstration density impact and validate H2R’s robustness, we retrain the diffusion policy with a dense RLBench dataset under identical settings. Table 5 shows higher data density significantly boosts overall success rate (e.g., MAE average from 3.3% to 45.0%). Critically, H2R consistently enhances performance across sparse and dense data regimes, confirming its general effectiveness regardless of demonstration density.
Table 5: Effects of Demonstration Density. Success rates (%) across imitation learning tasks in the RLBench environments with sparse data and dense data.
Data Model RLBench CloseBox CloseLaptopLid ToiletSeatDown Average Sparse data MAE 0.0 10.0 0.0 3.3 MAE+H2R 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) Dense data MAE 50.0 40.0 45.0 45.0 MAE+H2R 60.0 (+10.0%) 45.0 (+5.0%) 60.0 (+15.0%) 55.0 (+10.0%)
@@ 5 Real World Experiment
Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 23.3 43.3 (+20.0%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%) Franka-PlaceToy 10 15 (+5%) 10 25 (+15%) Franka-PlaceToy 25 40 (+15%) 20 30 (+10%) Franka-SweepRubbish 15 15 (0%) 20 20 (0%) Franka-SweepRubbish 20 35 (+15%) 20 20 (0%) Franka-WeightSauce 40 45 (+5%) 20 30 (+10%) Franka-WeightSauce 30 35 (+25%) 20 25 (+5%) Average 21.7 25 (+3.3%) 17.7 25 (+7.3%) Average 25 36.7 (+16.7%) 20 25 (+5%)
Robot Setups. We evaluate the effectiveness of H2R across three real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [RobotiqAdaptiveGrippers], (ii) a UR5 arm equipped with a Leaphand end effector [shaw2023leaphandlowcostefficient], and (iii) a dual-arm Franka Emika system using parallel Grippers. Realsense [IntelRealSenseL515] is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model. We provide the visualization of the robot setup in real-world scenarios in the supplementary materials.
Figure 4: Visualization of Real-world Manipulation Tasks. The left columns show Gripper tasks and the right columns show Leaphand tasks. Each task is illustrated with three frames, demonstrating the progression from the initial state to the completion of the manipulation.
Real-world Tasks. We set up nine tasks for gripper manipulation, dexterous manipulation and dual-arm manipulation (Figure 4). For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed. For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened. For the dual-arm manipulation tasks, we design: (7) Dualarm-PlaceToy, where the left arm takes a toy from the pink box and the right arm moves it to the blue box; (8) Dualarm-SweepRubbish, where the left arm sweeps the trash into the dustpan held by the right arm; and (9) Dualarm-WeightSauce, where the left arm places the cup on the scale and the right arm picks up the bottle for a pouring motion. We adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 8.
@@
Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800 Dualarm-PlaceToy 1000 300 1500 Dualarm-SweepRubbish 1000 300 1500 Dualarm-WeightSauce 1000 300 1500
Table 9: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
Policy Training. For policy training, we select the Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.